You are designing a pendulum clock to have a period of 4.7s. How long should the pendulum be?
1
Expert's answer
2020-03-04T10:02:58-0500

The period of oscillation of a pendulum


"T=2\pi\sqrt{\frac{l}{g}}."

The length of a pendulum


"l=\frac{gT^2}{4\pi^2}=\frac{9.8\times 4.7^2}{4\pi^2}=5.48\:\rm m."